Please can you help solve: 0.75x²-2.4x+75.Calculate the value of x when h = 73.08
Answer by stanbon(75887)

h = 0.75x²-2.4x+75.Calculate the value of x when h = 73.08
-----------------------------
0.75x²-2.4x+75= 73.08
0.75x^2-2.4x+1.92
Multiply thru by 100 to get rid of the decimals:
75x^2 - 240x +192 = 0
x = [240 +- sqrt(240^2-4*75*192)]/150
x = [240 +- 0]/150
x = 8/5
=0.16
==============
Cheers,
Stan H.